# Scalable Multi-Robot Motion Planning Using Guidance-Informed Hypergraphs

### v3

@@ V-D Kinodynamic Experiments
To evaluate Workspace Guided-DaSH with kinodynamic constraints, we use Kinodynamic Workspace Guided-DaSH (K-WG-DaSH) incorporating the modifications in Section IV-C. Kinodynamic-CBS [21] (K-CBS) is utilized as our baseline kinodynamic planner as it uses the same Kinodynamic RRT we employ in K-WG-DaSH, however they utilize a merge bound [26] creating composite robots when more coordination is needed. This will allow us to evaluate the benefits of the coordination guidance provided by K-WG-DaSH. Each method was given $45$ minutes to find a solution to planning problems in the warehouse environment, described in Section V-B1, using robots with $2$nd-order car dynamics.
@@
In the kinodynamic warehouse environment (results in Tab.  V), K-WG-DaSH finds solutions for problems with up to 6 robots consistently while K-CBS begins to struggle at 4 robots. In the 2 robot scenario, K-WG-DaSH maintains a high success rate with a lower planning time at the expense of a higher path cost as opposed to K-CBS which is inconsistent at finding solutions with a higher planning time but with a lower path cost. We attribute the scalability to the topological and coordination guidance provided by K-WG-DaSH.
@@
VI Conclusion and Future Work
In this work, we present both the Guided-DaSH framework for hybrid multi-robot planning and Workspace Guided-DaSH, an application of this framework to mobile robot MRMP in congested environments. Our adapted framework builds off of a state-of-the-art hybrid planning framework, DaSH, using guidance to enable its use for a wider range of multi-robot planning problems which lack a highly structured search space. Additionally, we add support for kinodynamic planning via a modified conflict resolution structure, which also boosts the scalability of our framework by eliminating unnecessary work when finding motion solutions.
Workspace Guided-DaSH leverages topological guidance in multiple ways, both to move between different levels of composition during planning and to guide sampling through narrow passages in the workspace. By using knowledge of the workspace, our method is able to efficiently find paths for robot groups of size up to an order of magnitude larger than existing state-of-the-art methods. Building off of this work, we plan to explore the use of different kinodynamic low-level motion planners which plan more intelligently for multi-robot systems by interleaving sampling-based planning and trajectory optimization [27]. Additionally, we plan to explore using parallel computing to further improve the performance of this method, taking advantage of the hypergraph’s ability to naturally decompose the multi-robot planning problems.

### v4

@@ IV-D Kinodynamic Experiments
To evaluate Workspace Guided-DaSH with kinodynamic constraints, we use Kinodynamic Workspace Guided-DaSH (K-WG-DaSH) incorporating the modifications in Sec. III-D. We compare our approach to Kinodynamic-CBS [kal-cbsfmrmpwkc-22] (K-CBS) and Kinodynamic-ARC [qsmma-karcfmrkp-2025] (K-ARC). K-CBS uses the same Kinodynamic RRT we employ in K-WG-DaSH, however, they utilize a merge bound [ssfs-macbsfomapp-12] creating composite robots when more coordination is needed. K-ARC uses the same reactive coordination approach employed by ARC, but additionally uses trajectory optimization to ensure kinodynamic feasibility. These allow us to evaluate the benefits of the guidance provided by K-WG-DaSH. Each method was given $45$ minutes to find a solution to planning problems in the warehouse environment, described in Section IV-B1, using robots with 22nd-order car dynamics:
$\dot{x}=v\cos{\theta},\ \dot{y}=v\sin{\theta},\ \dot{\theta}=\frac{v}{l}\tan{\phi},\ \dot{v}=a,\ \dot{\phi}=\gamma\$(1)
@@
Results for the kinodynamic warehouse environment are shown in Tab. V. K-ARC was unable to solve the 2-robot scenario within the time limit and is omitted from Tab. V. K-ARC struggles in this scenario because of both the large size of the environment and the constrained space in which both robots must move, which make the trajectory optimization problem difficult to solve. K-WG-DaSH finds solutions for problems with up to 6 robots consistently while K-CBS begins to struggle at 4 robots. In the 2-robot scenario, K-WG-DaSH maintains a high success rate with a lower planning time at the expense of a higher path cost as opposed to K-CBS which is inconsistent at finding solutions with a higher planning time but with a lower path cost. We attribute the scalability to the topological guidance provided by K-WG-DaSH.
@@
V Conclusion and Future Work
In this work, we present Workspace Guided-DaSH, a novel hybrid MRMP method for environments with narrow passages. Our method builds off of a state-of-the-art hybrid planning framework, DaSH, using workspace guidance to enable its use for a multi-robot planning problem that lacks a highly structured search space. Additionally, we add support for kinodynamic planning via a modified conflict resolution structure, which also boosts the scalability of our method by eliminating unnecessary work when finding motion solutions.
Workspace Guided-DaSH leverages topological guidance in multiple ways, both to move between different levels of composition during planning and to guide sampling through narrow passages in the workspace. By using knowledge of the workspace, our method is able to efficiently find paths for robot groups of size up to an order of magnitude larger than existing state-of-the-art methods. Building off of this work, we plan to explore the use of different kinodynamic low-level motion planners which plan more intelligently for multi-robot systems by interleaving sampling-based planning and trajectory optimization [moth-ddbcbsfmrkmp-23]. Additionally, we plan to explore using parallel computing to further improve the performance of this method, taking advantage of the hypergraph’s ability to naturally decompose the multi-robot planning problems.